# TrackCore-F: Deploying Transformer-Based Subatomic Particle Tracking on FPGAs

### v1

@@ Abstract
The Transformer Machine Learning (ML) architecture has been gaining considerable momentum in recent years. In particular, computational High-Energy Physics tasks such as jet tagging and particle track reconstruction (tracking), have either achieved proper solutions, or reached considerable milestones using Transformers. On the other hand, the use of specialised hardware accelerators, especially FPGAs, is an effective method to achieve online, or pseudo-online latencies. The development and integration of Transformer-based ML to FPGAs is still ongoing and the support from current tools is very limited to non-existent. Additionally, FPGA resources present a significant constraint. Considering the model size alone, while smaller models can be deployed directly, larger models are to be partitioned in a meaningful and ideally, automated way. We aim to develop methodologies and tools for monolithic, or partitioned Transformer synthesis, specifically targeting inference. Our primary use-case involves two machine learning model designs for tracking, derived from the TrackFormers project. We elaborate our development approach, present preliminary results, and provide comparisons.
@@ 1 Introduction
The move to ML-assisted tracking solutions is deemed necessary and inevitable. While design efforts are moving forward, another important aspect to consider is deployment. The accurate tracking has been a post-mortem task, which is due to the computation requirements. While the introduction of ML algorithms will improve computational performance, the increase in scale and frequency of experiments will somewhat counter and reduce effects of such improvements. This is especially the case for the scales expected from the upcoming High-Luminosity stage of the LHC (HL-LHC). On top of that, efficient execution of ML algorithms has predominantly been tied to GPUs as the platform of choice. While not a universal dependency, GPUs dominate the ML deployment scene. Having said that, there are viable alternative forms of hardware acceleration, e.g., FPGA, custom ASIC, and Neuromorphic. We aim to deploy ML-assisted track reconstruction on FPGAs to achieve better or equivalent latencies. FPGAs enable on-site deployment of tracking and entail considerable energy efficiency potential.
@@
Deployment of models based on the Transformer architecture on hardware, especially FPGAs, has been a point of interest for the research community. This is especially noticeable during the past and the current year, 2024–2025. Focusing on the optimisations, 4 main trends are observable, namely: approaches focusing on quantization techniques [1, 2, 3], approaches utilising off-chip memory [4], approaches levering sparsity trough pruning [5, 6, 7], and approaches based on optimisation of non-linear functions such as SoftMax [8].
One of the important metrics to consider when evaluating implementations is throughput. At the time of this writing, the most promising approaches demonstrating the highest throughput are [1] and [9], with the former achieving noticeably higher throughput.
@@ 2 Background
At the Large Hadron Collider (LHC), particles are accelerated in opposite directions in a circular accelerator and made to collide at four interaction points, where large-scale detectors are positioned. These major detectors are ALICE [10], ATLAS [11], CMS [12], and LHCb [13]. Detectors perform two key functions: tracking and calorimetry, allowing the calculation of their momentum, $p$ and measuring the energy, $E$, deposited by particles, respectively. Together, these measurements enable the calculation of a particle’s mass, $m$, using the relativistic energy-momentum relation: $E^{2}=(mc^{2})^{2}+(pc)^{2}$, where $c$ is the speed of light. Accurately determining particle mass is essential for identifying known particles and discovering new ones.
@@ 2.1 Tracking algorithms
There has been continuous efforts channelled into the design and development of ML-based, or rather ML-assisted, tracking solutions. Two ML model architectures stand out: Graph Neural Networks (GNNs) and more recently, Transformers [14]. Within the scope of this paper, we focus on two Transformer designs from the project TrackFormers [15], EncCla and EncReg. Both models operate as so-called single-shot models, i.e., they take in a full event’s data and perform hit to track association for the whole event.
The Encoder-Classifier (EncCla) is an encoder-only Transformer design. This approach takes in the coordinates from an event and predicts their association to pre-defined class labels. Class labels, ensured to be unique, are generated through binning of the track parameter space. The largest variant has close to 1.5 million parameters with estimated memory consumptions of 5.69 MB and 0.07 MB for parameters and activations, respectively.
The Encoder-Regressor (EncReg) is also an encoder-only Transformer design. It does not rely on class labels, but regression of potential tracks’ parameters for a single event. As a post-processing step, a clustering algorithm has to be applied to model’s output. As such, predicted track parameters per hit are clustered, forming track associations. EncReg uses HDBSCAN to achieve this clustering. The model has close to 76 484 parameters with estimated memory consumptions of 0.29 MB and 0.07 MB for parameters and activations, respectively.
@@ 2.2 Datasets
The two model designs have been trained with 5 different datasets, forming a progression of simple to complex representations of tracks and hits, i.e., track function complexity, track count, and by extension, hit count: 10–50 (variable count) linear tracks per event (REDVID), 10–50 (variable count) helical tracks per event (REDVID), 50–100 (variable count) helical tracks per event (REDVID), 10–50 (variable count) tracks per event (TrackML), 200–500 (variable count) tracks per event (TrackML). The first three datasets are the result of simulations using REDVID simulation framework [16]. The last two datasets on are scale-reduced versions of the data associated with the TrackML Kaggle challenge [17].
@@ 3 Implementation and results
The utilised test bench is an ARM Zynq UltraScale+ MPSoC ZCU102 evaluation kit. The on-board EG device (ZU9EG) consists of a Quad-Core ARM Cortex-A53 Processing Unit (PU) and a Programmable Logic (PL) with the following specification [18]: System Logic Cells 599 550, Configurable Logic Block (CLB) Flip-Flops 548 160, CLB LUTs 274 080, Distributed RAM (Mb) 8.8, Block RAM (Mb) 32.1, DSP Slices 2 520.
A variety of tooling has to be used in a progression to achieve a deployable synthesised kernel. In particular, for pre-trained ML models, we have taken advantage of the following:
  • PyTorch: Our models are provided in the PyTorch format, which could optionally be used for quantization.
  • ONNX: Open Neural Network Exchange (ONNX) is an open-source format for representing ML models, which reveals low-level operations, input/output dimensions, data types, and weight Tensor values (if present) per operation. We have used the ONNX format for quantization as well.
@@
  • AMD Vivado 2022.2: Vivado Design Suite is used to integrate and configure the IP designed using Vitis HLS with the Zynq MPSoC and synthesise the final bitstream.
@@ 3.1 Development flow
As depicted in Figure˜1, our deployment workflow is comprised of different steps involving the aforementioned tooling.
@@
Focusing on a single encoder layer, the model is split into three parts: the segment before the first encoder layer, the encoder layer to be implemented and deployed on the FPGA, and the remaining model segment. The encoder kernel block is developed using HLS C/C++ in Vitis. The top-level function takes all the weights of the encoder layer as input. These weights are transferred to the kernel using a AXI4-Lite interface. Once all the parameters are loaded in, the computations for the encoder layer take place. This consists of the Multi-Head Attention (MHA), followed by the first Add AND Normalisation, a feed-forward layer, and finally a last Add AND Normalisation layer. HLS Pragmas are used to optimise performance.
@@ 3.3 Resource consumption
Table 2: ZCU102 FPGA resource utilisation for a single encoder layer.
@@
As it can be seen, the limiting factor is the available BRAM, which has a utilisation of 76.1%. However, the number of used BRAMs can be reduced effectively by considering more DDR memory at the expense of increased memory transfer and latency. Thus, given that the BRAM utilisation can be reduced, the second limiting factor will be the number of LUTs, which are at 24.01% utilisation for a single encoder layer. This suggest that a maximum of 4 encoder layers can be implemented on this particular FPGA and for this model.
@@ 4 Conclusion
We provided a structured development flow for deployment of pre-trained tracking models on FPGAs. A partial deployment is as feasible as a full deployment. While not as performant, we see a partial deployment as valuable as a full one, since it enables the deployment of model inference on more accessible hardware. We notice that application of optimisations could be more important than the HLS implementation itself. We have seen how detrimental quantization can be for the model’s prediction accuracy. Optimisations and their effects are highly dependent on the model itself and the HLS implementations.

### v2

@@ Abstract
The Transformer Machine Learning (ML) architecture has been gaining considerable momentum in recent years. In particular, computational High-Energy Physics tasks such as jet tagging and particle track reconstruction (tracking), have either achieved proper solutions, or reached considerable milestones using Transformers. On the other hand, the use of specialised hardware accelerators, especially FPGAs, is an effective method to achieve online, or pseudo-online latencies. The development and integration of Transformer-based ML to FPGAs is still ongoing and the support from current tools is very limited or non-existent. Additionally, FPGA resources present a significant constraint. Considering the model size alone, while smaller models can be deployed directly, larger models are to be partitioned in a meaningful and ideally, automated way. We aim to develop methodologies and tools for monolithic, or partitioned Transformer synthesis, specifically targeting inference. Our primary use-case involves two machine learning model designs for tracking, derived from the TrackFormers project. We elaborate our development approach, present preliminary results, and provide comparisons.
@@ 1 Introduction
The move to ML-assisted tracking solutions is deemed necessary and inevitable. While design efforts are moving forward, another important aspect to consider is deployment. The accurate tracking has been a post-mortem task, which is due to the computational requirements. While the introduction of ML algorithms will improve the computational performance, the increase in scale and frequency of experiments will somewhat counter and reduce effects of such improvements. This is especially the case for the scales expected from the upcoming High-Luminosity stage of the LHC (HL-LHC). On top of that, efficient execution of ML algorithms has predominantly been tied to Graphics Processing Units (GPUs) as the platform of choice. While not a universal dependency, GPUs dominate the ML deployment scene. Nevertheless, there are viable alternative forms of hardware acceleration, e.g., Field-Programmable Gate Array (FPGA), custom Application-Specific Integrated Circuit (ASIC), and Neuromorphic. We aim to deploy ML-assisted track reconstruction on FPGAs to achieve better or equivalent latencies. Additionally, FPGAs enable on-site deployment of tracking and entail considerable energy efficiency potential.
@@
Deployment of models based on the Transformer architecture on hardware, especially FPGAs, has been a point of interest for the research community. This is especially noticeable during the past and the current year, 2024–2025. Focusing on the optimisations, 4 main trends are observable, namely: approaches focusing on quantization techniques [Guo:2025:VTAH, Zhao:2025:EVTA, Du:2024:EEFB], approaches utilising off-chip memory [Kabir:2025:RATN], approaches leveraging sparsity through pruning [Zhang:2024:EFTA, Wang:2024:EFTA, Li:2024:EEFA], and approaches based on optimisation of non-linear functions such as SoftMax [Bai:2024:SEWA].
One of the important metrics to consider when evaluating implementations is throughput. At the time of this writing, the most promising approaches demonstrating the highest throughput are [Guo:2025:VTAH] and [He:2025:FTFA], with the former achieving noticeably higher throughput.
@@ 2 Background
At the Large Hadron Collider (LHC), particles are accelerated in opposite directions in a circular accelerator and made to collide at four interaction points, where large-scale detectors are positioned. These major detectors are ALICE [Collaboration:2008:ALICE], ATLAS [Collaboration:2008:ATLAS], CMS [Collaboration:2008:CMS], and LHCb [Collaboration:2008:LHCb]. Detectors perform two key functions: tracking and calorimetry, allowing the calculation of particle momentum, $p$ and measuring the energy, $E$, deposited by particles, respectively. Together, these measurements enable the calculation of a particle’s mass, $m$, using the relativistic energy-momentum relation: $E^{2}=(mc^{2})^{2}+(pc)^{2}$, where $c$ is the speed of light. Accurately determining particle mass is essential for identifying known particles and discovering new ones.
@@ 2.1 Tracking algorithms
There has been continuous efforts channelled into the design and development of ML-based, or rather ML-assisted, tracking solutions. Two ML model architectures stand out: Graph Neural Networks (GNNs) and more recently, Transformers [Vaswani:2017:Attention]. Within the scope of this paper, we focus on two Transformer designs from the project TrackFormers [Caron:2025:TrackFormers], EncCla and EncReg. Both models operate as so-called single-shot models, i.e., they take in a full event’s data and perform hit to track association for the whole event.
The Encoder-Classifier (EncCla) is an encoder-only Transformer design. This approach takes in hit coordinates from an event and predicts their association to pre-defined class labels. Class labels, ensured to be unique, are generated through binning of the track parameter space. The largest variant has close to 1.5 million parameters with estimated memory consumptions of 5.69 MB and 0.07 MB for parameters and activations, respectively.
The Encoder-Regressor (EncReg) is also an encoder-only Transformer design. It does not rely on class labels, but instead regresses the parameters of potential tracks for a single event. As a post-processing step, a clustering algorithm has to be applied to model’s output. As such, predicted track parameters per hit are clustered, forming track associations. EncReg uses HDBSCAN to achieve this clustering. The model has close to 76 484 parameters with estimated memory consumptions of 0.29 MB and 0.07 MB for parameters and activations, respectively.
@@ 2.2 Datasets
The two model designs have been trained with 5 different datasets, forming a progression of simple to complex representations of tracks and hits, i.e., track function complexity, track count, and by extension, hit count: 10–50 (variable count) linear tracks per event (REDVID), 10–50 (variable count) helical tracks per event (REDVID), 50–100 (variable count) helical tracks per event (REDVID), 10–50 (variable count) tracks per event (TrackML), 200–500 (variable count) tracks per event (TrackML). The first three datasets are the result of simulations using REDVID simulation framework [Odyurt:2024:RSHE]. The last two datasets are scale-reduced versions of the data associated with the TrackML Kaggle challenge [Kiehn:2019:TrackML].
@@ 3 Implementation and results
The utilised test bench is an ARM Zynq UltraScale+ MPSoC ZCU102 evaluation kit. The on-board EG device (ZU9EG) consists of a Quad-Core ARM Cortex-A53 Processing Unit (PU) and a Programmable Logic (PL) with the following specification [AMD:2025:Zynq]: System Logic Cells 599 550, Configurable Logic Block (CLB) Flip-Flops 548 160, CLB LUTs 274 080, Distributed RAM (Mb) 8.8, Block RAM (Mb) 32.1, DSP Slices 2 520.
A variety of tools must be used in sequence to obtain a deployable synthesised kernel. In particular, for pre-trained ML models, the following have been used:
  • PyTorch: Models are provided in the PyTorch format, which could optionally be used for quantization.
  • ONNX: Open Neural Network Exchange (ONNX) is an open-source format for representing ML models, which reveals low-level operations, input/output dimensions, data types, and weight Tensor values (if present) per operation. The ONNX format is utilised for quantization as well.
@@
  • AMD Vivado 2022.2: Vivado Design Suite is used to integrate and configure the IP designed using Vitis HLS with the Zynq MPSoC and to synthesise the final bitstream.
@@ 3.1 Development flow
As depicted in Figure˜1, our deployment workflow is comprised of different steps involving the aforementioned tools.
@@
Focusing on a single encoder layer, the model is split into three parts: the segment before the first encoder layer, the encoder layer to be implemented and deployed on the FPGA, and the remaining model segment. The encoder kernel block is developed using HLS C/C++ in Vitis. The top-level function takes all the weights of the encoder layer as input. These weights are transferred to the kernel using an AXI4-Lite interface. Once all the parameters are loaded in, the computations for the encoder layer take place. This consists of the Multi-Head Attention (MHA), followed by the first Add AND Normalisation, a feed-forward layer, and finally a last Add AND Normalisation layer. HLS Pragmas are used to optimise performance.
@@ 3.3 Resource consumption
Table 2: ZCU102 FPGA resource utilisation for a single encoder layer, as reported by Vitis.
@@
As it can be seen, the limiting factor is the available BRAM, which has a utilisation of 38.04%. However, the number of used BRAMs can be reduced effectively by considering more DDR memory at the expense of increased memory transfer and latency. Thus, given that the BRAM utilisation can be reduced, the second limiting factor will be the number of LUTs, which is at 24.01% utilisation for a single encoder layer. This suggests that a maximum of 4 encoder layers can be implemented on this particular FPGA and for this model.
@@ 4 Conclusion
We have presented a structured development flow for the deployment of pre-trained tracking models on FPGAs. A partial deployment is as feasible as a full deployment. Although not as performant, we consider a partial deployment as valuable as a full one, since it enables model inference to be deployed on more accessible hardware. We observe that the application of optimisations can be more important than the HLS implementation itself. We have also shown how detrimental quantization can be to the model’s prediction accuracy. Optimisations and their effects are highly dependent on the model and the HLS implementation.